OpenStudy (anonymous):

Which one is bigger, 5 3/4 or 5.825???

OpenStudy (anonymous):

So remember that with a mixed fraction, you can convert the fractional part to a decimal and just add. 3/4 as a decimal is 0.75, and 5+0.75=5.75. Should be easier to compare now :)

OpenStudy (anonymous):

Can you explain this on a 5th grader's level?

OpenStudy (anonymous):

It is due Monday!

OpenStudy (anonymous):

To compare the two kinds of numbers (mixed fraction and decimal), it's easier to have them both be in the same form. One way to do it in this case is to convert \(5\frac{3}{4}\) to a decimal. \(5\frac{3}{4}\) is the same as \(5+\frac{3}{4}\). Since 5 is already a decimal, all you have to do is convert \(\frac{3}{4}\) into a decimal. You can do this by plugging it in your calculator and finding the result is 0.75. Adding 5 and 0.75 gives you 5.75, so the question now becomes, is 5.75 bigger than 5.825, or not?

OpenStudy (anonymous):

Oh!!! So 5.75 is bigger?

OpenStudy (anonymous):

Well... what happens if you subtract 5.825 from 5.75?

OpenStudy (anonymous):

Ummm, it is o.75?

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Not quite. Let's take a step back. What happens if you subtract five from 5.75? Remember that 5 is the same as 5.00 (the zeros after the decimal don't matter), so it's the same as: \[ \begin{align*} 5&.75 -5&.00 \end{align*} \]

OpenStudy (anonymous):

\[ \begin{align} 5&.75\\ -5&.00 \end{align} \]

OpenStudy (anonymous):

75?

OpenStudy (anonymous):

Don't forget to drop down the decimal point!

OpenStudy (anonymous):

so .75?

OpenStudy (anonymous):

Exactly. Now what happens if you subtract 5 from 5.825?

OpenStudy (anonymous):

If you're having trouble, do the same thing I did above. This would look like: \[ \begin{align} 5&.825\\ -5&.000 \end{align} \] Since the 0s don't matter again.

OpenStudy (anonymous):

825?

OpenStudy (anonymous):

The poor decimal point got left out again :)

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

What do I do now?

OpenStudy (anonymous):

Now you've got 0.825 and 0.75. Can you see which of those is bigger?

OpenStudy (anonymous):

Yes. 0.825!

OpenStudy (anonymous):

Excellent! Now, you subtracted 5 from both 5.825 and from 5.75 to get 0.825 and 0.75. So since 0.825 is bigger than 0.75, what does that say about which one is bigger between 5.825 and 5.75?

OpenStudy (anonymous):

5.825 is bigger than 5 3/4?

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Yes! Sorry, went away for a bit.

OpenStudy (anonymous):

Can you help me some more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

How do add fractions?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Okay, to add fractions you need to make sure they have the same denominator. Can you give me two example ones you need to add?

OpenStudy (anonymous):

1/2 and 5/9? I have to write it as a mixed number in simplest form.

OpenStudy (anonymous):

You are a great toutor :)

OpenStudy (anonymous):

Okay. So like I said, the first step is making sure they have the same denominator. To do that, first you have to understand that if I multiply a fraction by a number over itself, then it stays the same. So, as an example, if I were to multiply a fraction by \(\frac{3}{3}\) or \(\frac{5}{5}\) or \(\frac{9}{9}\), then it would stay the same. Are you clear with that?

OpenStudy (anonymous):

Nope! I am only in 5th grade.

OpenStudy (anonymous):

So what do you get if you divide two by two?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Exactly. And if you divide three by three, or four by four, or 17 by 17, you still get one, right?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

The next step....

OpenStudy (anonymous):

Awesome. And what happens if you multiply a number by one?

OpenStudy (anonymous):

The same number.

OpenStudy (anonymous):

Does this have something to do with fractions?

OpenStudy (anonymous):

Sorry, that was rude :(

OpenStudy (anonymous):

Can you please contuine?

OpenStudy (anonymous):

I'm sorry, my Internet died :/ If you still want help, I'm willing to continue now that it's back up.